-x-(13+4x)=-3(5-9x)+2
Answer by unlockmath(1688)

Hello,
For this one, -x-(13+4x)=-3(5-9x)+2 let's expand it out to:
-x-13-4x=-15+27x+2
Combine like terms:
-5x-13=27x-13
Add 13 and 5x to get:
33x=0
x=0
Plug in this to see if it is correct.
RJ
www.math-unlock.com